OpenStudy (anonymous):

find the center and radius of the circle x squared -5y+y squared +4x-1 =0

OpenStudy (anonymous):

center = (-2,5/2) radius=sqrt(45)/2

OpenStudy (anonymous):

how did you do it chanh

OpenStudy (anonymous):

x^2+4x+y^2-5y-1=0 x^2+4x+4+y^2-5y+25/4-4-25/4-1=0 (x+2)^2+(y-5/2)^2=45/4 thus center= (-2,5/2) \[radius=\sqrt{45}/2\]

OpenStudy (anonymous):

thank you chanh

OpenStudy (anonymous):

u welcome